### v1

@@ 2.1 Instantaneous Soundness Checking
For all our benchmarks, we use the hyperfine benchmarking tool [11] (version 1.18.0), which calculates the average runtime when executing each soundness check ten or more times. The benchmarks were run on Ubuntu 22.04.4 with an AMD Ryzen 7700X processor (4.5GHz) and 32 GB of RAM. All used BPMN models, our tools to generate them, and benchmarking scripts are given in the artifacts of this paper [12].
@@ 3.1 Tool overview
We developed soundness checking in Rust due to its memory efficiency, absence of garbage collection and lightweight hardware abstractions, resulting in an execution performance comparable to C or C++, while retaining memory safety guarantees. A fast programming language and direct implementation of the BPMN semantics play the key role in achieving instantaneous soundness checking. We would like to emphasize that the current implementation does feature specific optimization techniques such as partial order reduction, therefore, we assume that performance may be amplified even further for certain types of models or violations, see, e.g., the parallel branches example.
@@ 3.2 Soundness checking in Rust
Similarly, for Option To Complete, we examine whether a state is stuck. This means that some process instance is still ongoing, indicated by the presence of tokens, and the state has no outgoing transitions. In contrast, we check No Dead Activities by remembering all executed activities during state space generation and comparing them to all activities in the model. Furthermore, Proper Completion is checked by remembering executed end events in each state.
@@ 4 Related Work
BPMN specification coverage: Most related work focuses on the depth of BPMN formalization while providing soundness and safeness checking. Thus, these approaches show how different BPMN elements can be formalized and compare themselves with each other regarding supported elements [4, 8, 9, 10]. The supported BPMN elements come close to the capabilities of popular process orchestration platforms. Our tool supports the same depth of BPMN elements as [4], i.e., the most used gateways, tasks, and events. A detailed comparison is available in [12], and we plan to support more elements in the future. We focus on tool performance and capabilities concerning soundness comprehension, soundness resolution, and seamless integration into BPMN modeling tools.
@@
Tool capabilities: Another way to compare the different BPMN formalizations and soundness-checking tools is to investigate their capabilities. Most tools formalize large parts of the BPMN specification and allow soundness and safeness checking [4, 8, 9, 10]. Some tools investigate additional aspects such as the introduction of data and time during verification [5, 8]. Furthermore, other tools allow specifying and checking custom temporal logic properties [4, 9] and even provide graphical interfaces to easy the specification [10] Moreover, some tools provide interactive BPMN simulation using token-flow animation [1, 5], while others visualize counterexamples for soundness violations using tokens [8].
Our tool focuses on instantaneous soundness and safeness checking and does not support custom properties or data and time. We do not provide BPMN simulation since other tools already offer this. However, we use token-flow animation to visualize counterexamples for soundness violations interactively. This improves comprehension compared to previous static, less interactive visualizations. In addition, to the best of our knowledge, our tool is the only one that provides quick fixes, i.e., automatic resolutions if soundness properties are violated. To sum up, our tool incorporates and enhances several ideas from the state of the art while adding novel concepts, such as quick fixes. We advocate for a pragmatic approach, prioritizing performance and understanding above all else to ensure seamless integration into BPMN modeling tools.

### v2

@@ 2.1 Instantaneous Soundness Checking
For all our benchmarks, we use the hyperfine benchmarking tool [11] (version 1.18.0), which calculates the average runtime when executing each soundness check ten or more times. The benchmarks were run on Ubuntu 22.04.4 with an AMD Ryzen 7700X processor (4.5GHz) and 32 GB of RAM (5600 MHz). All used BPMN models, our tools to generate them, and benchmarking scripts are given in the artifacts of this paper [12].
@@ 3.1 Tool overview
We developed soundness checking in Rust due to its memory efficiency, absence of garbage collection and lightweight hardware abstractions, resulting in an execution performance comparable to C or C++, while retaining memory safety guarantees. A fast programming language and direct implementation of the BPMN semantics are key in achieving instantaneous soundness checking. We would like to emphasize that the current implementation does not include optimization techniques such as partial order reduction, therefore, we assume that performance may be amplified even further for certain types of models or violations, see, e.g., the parallel branches example.
@@ 3.2 Soundness checking in Rust
Similarly, we examine whether a state is stuck for Option To Complete. This means that some process instances are still ongoing, indicated by the presence of tokens, and the state has no outgoing transitions. In contrast, we check No Dead Activities by remembering all executed activities during state space generation and comparing them to all activities in the model. Furthermore, Proper Completion is checked by remembering executed end events in each state.
@@ 4 Related Work
BPMN specification coverage: Most related work focuses on the depth of BPMN formalization while providing soundness and safeness checking. Thus, these approaches show how different BPMN elements can be formalized and compare themselves with each other regarding supported elements [4, 8, 9, 10]. The supported BPMN elements come close to the capabilities of popular process orchestration platforms. Our tool supports the same depth of BPMN elements as [4], i.e., the most used gateways, tasks, and events. A detailed comparison is shown in Table 3, and we plan to support more elements in the future [12]. We focus on tool performance and capabilities concerning soundness comprehension, soundness resolution, and seamless integration into BPMN modeling tools.
Table 3: BPMN elements supported by different formalizations (based on [14]).
BPMN element/feature Kräuter [9] Corradini [4] Houhou [8] This paper
Instantiation and termination
Start event instantiation X X X X
Exclusive event-based gateway X
   instantiation
Parallel event-based gateway
   instantiation
Receive task instantiation X X
Normal process completion X X X X
Activities
Activity X X X X
Subprocess X X
Ad-hoc subprocesses
Loop activity
Multiple instance activity
Transaction
Gateways
Parallel gateway X X X X
Exclusive gateway X X X X
Inclusive gateway (split) X X X
Inclusive gateway (merge) X X
Event-based gateway X X X X
Complex gateway
Events
None Events X X X X
Message events X X X X
Timer Events X
Escalation Events X
Error Events X
Cancel Events
Compensation Events
Conditional Events
Link Events X X
Signal Events X
Multiple Events
Terminate Events X X X X
Boundary Events X X
Event subprocess X
@@
Tool capabilities: Another way to compare the different BPMN formalizations and soundness-checking tools is to investigate their capabilities. Most tools formalize large parts of the BPMN specification and allow soundness and safeness checking [4, 8, 9, 10]. Some tools investigate additional aspects such as the introduction of data and time during verification [5, 8]. Furthermore, other tools allow specifying and checking custom temporal logic properties [4, 9] and even provide graphical interfaces to ease the specification [10] Moreover, some tools provide interactive BPMN simulation using token-flow animation [1, 5], while others visualize counterexamples for soundness violations using tokens [8].
Our tool focuses on instantaneous soundness and safeness checking and does not support custom properties, data or time. We do not provide BPMN simulation since other tools already offer this. However, we use token-flow animation to visualize counterexamples for soundness violations interactively. This improves comprehension compared to previous static, less interactive visualizations. In addition, to the best of our knowledge, our tool is the only one that provides quick fixes, i.e., automatic resolutions if soundness properties are violated. To sum up, our tool incorporates and enhances several ideas from the state of the art while adding novel concepts, such as quick fixes. We advocate for a pragmatic approach, prioritizing performance and understanding above all else to ensure seamless integration into BPMN modeling tools.